# Binary superfluids: low-energy properties and dissipative processes from spontaneous emission of massive phonons

### v1

@@ A.1 Expansion around mean field of a generic Lagrangian
where $\rho$ is the radial field and $\theta$ is the cyclic coordinate. We assume vanishing temperature and that the Lagrangian is invariant under a global $U(1)$ transformation. In the ground-state, such group is assumed to be spontaneously broken.
@@
${\cal L}={\cal L}_{0}(\rho,\partial_{\mu}\rho,\partial_{\mu}\theta)+\tilde{\rho}\left.\left(\frac{\delta{\cal L}}{\delta\rho}-\partial_{\mu}\frac{\delta{\cal L}}{\delta\partial_{\mu}\rho}\right)\right|_{\rho,\theta}+\delta\partial_{\mu}\tilde{\theta}\left.\frac{\delta{\cal L}}{\delta\partial_{\mu}\theta}\right|_{\rho,\theta}+{\cal L}_{2}+\dots$ (165)
@@
${\cal L}_{2}=$ $\frac{\tilde{\rho}^{2}}{2}\left.\frac{\delta^{2}{\cal L}}{\delta\rho^{2}}\right|_{\rho,\theta}+\tilde{\rho}\partial_{\mu}\tilde{\rho}\left.\frac{\delta^{2}{\cal L}}{\delta\rho\delta\partial_{\mu}\rho}\right|_{\rho,\theta}+\frac{\partial_{\mu}\tilde{\rho}\partial_{\nu}\tilde{\rho}}{2}\left.\frac{\delta^{2}{\cal L}}{\delta\partial_{\mu}\rho\delta\partial_{\nu}\rho}\right|_{\rho,\theta}+\tilde{\rho}\partial_{\mu}\tilde{\theta}\left.\frac{\delta^{2}{\cal L}}{\delta\rho\delta\partial_{\mu}\theta}\right|_{\rho,\theta}\,+$
$+\partial_{\nu}\tilde{\rho}\partial_{\mu}\tilde{\theta}\left.\frac{\delta^{2}{\cal L}}{\delta\partial_{\nu}\rho\delta\partial_{\mu}\theta}\right|_{\rho,\theta}+\frac{\partial_{\nu}\tilde{\theta}\partial_{\mu}\tilde{\theta}}{2}\left.\frac{\delta^{2}{\cal L}}{\delta\partial_{\nu}\theta\delta\partial_{\mu}\theta}\right|_{\rho,\theta}\,,$ (166)
@@
We now discuss in some details the various terms in the right-hand side of Eq. (A.1). The first term is proportional to minus the curvature of the potential at the minimum:
@@
${\cal L}_{2}=$ $\frac{1}{2}\left(V^{\mu}\partial_{\mu}\tilde{\theta}-A\square\tilde{\theta}\right)\frac{1}{\square+\tilde{m}^{2}}\left(V^{\mu}\partial_{\mu}\tilde{\theta}-A\square\tilde{\theta}\right)+\frac{B^{2}\eta^{\mu\nu}}{2}\partial_{\nu}\tilde{\theta}\partial_{\mu}\tilde{\theta}\,,$ (175)
@@ A.2 Derivative expansion of the pressure
${\cal L}=P(\mu)-\frac{\partial P}{\partial\mu}v_{\mu}\partial^{\mu}\tilde{\theta}+\frac{1}{2}\left[\eta^{\mu\nu}\frac{1}{\mu}\frac{\partial P}{\partial\mu}+v^{\mu}v^{\nu}\left(\frac{\partial^{2}P}{\partial\mu^{2}}-\frac{1}{\mu}\frac{\partial P}{\partial\mu}\right)\right]\partial_{\nu}\tilde{\theta}\partial_{\mu}\tilde{\theta}\,,$ (180)

### v2

@@ A.1 Expansion around mean field of a generic Lagrangian
where $\rho$ is the radial field and $\theta$ is the cyclic coordinate. We assume vanishing temperature and that the Lagrangian is invariant under a global $U(1)$ transformation. In the ground state, such a group is assumed to be spontaneously broken.
@@
$+\delta\partial_{\mu}\tilde{\theta}\left.\frac{\delta{\cal L}}{\delta\partial_{\mu}\theta}\right|_{\rho,\theta}+{\cal L}_{2}+\dots$ (169)
@@
${\cal L}_{2}=$ $\frac{\tilde{\rho}^{2}}{2}\left.\frac{\delta^{2}{\cal L}}{\delta\rho^{2}}\right|_{\rho,\theta}+\tilde{\rho}\partial_{\mu}\tilde{\rho}\left.\frac{\delta^{2}{\cal L}}{\delta\rho\delta\partial_{\mu}\rho}\right|_{\rho,\theta}+\frac{\partial_{\mu}\tilde{\rho}\partial_{\nu}\tilde{\rho}}{2}\left.\frac{\delta^{2}{\cal L}}{\delta\partial_{\mu}\rho\delta\partial_{\nu}\rho}\right|_{\rho,\theta}$
$+\tilde{\rho}\partial_{\mu}\tilde{\theta}\left.\frac{\delta^{2}{\cal L}}{\delta\rho\delta\partial_{\mu}\theta}\right|_{\rho,\theta}\,+\partial_{\nu}\tilde{\rho}\partial_{\mu}\tilde{\theta}\left.\frac{\delta^{2}{\cal L}}{\delta\partial_{\nu}\rho\delta\partial_{\mu}\theta}\right|_{\rho,\theta}$
$+\frac{\partial_{\nu}\tilde{\theta}\partial_{\mu}\tilde{\theta}}{2}\left.\frac{\delta^{2}{\cal L}}{\delta\partial_{\nu}\theta\delta\partial_{\mu}\theta}\right|_{\rho,\theta}\,,$ (170)
@@
Let us discuss in some details the various terms in the right-hand side of Eq. (A.1). The first term is proportional to minus the curvature of the potential at the minimum:
@@
${\cal L}_{2}=$ $\frac{1}{2}\left(V^{\mu}\partial_{\mu}\tilde{\theta}-A\square\tilde{\theta}\right)\frac{1}{\square+\tilde{m}^{2}}\left(V^{\mu}\partial_{\mu}\tilde{\theta}-A\square\tilde{\theta}\right)$
$+\frac{B^{2}\eta^{\mu\nu}}{2}\partial_{\nu}\tilde{\theta}\partial_{\mu}\tilde{\theta}\,,$ (179)
@@ A.2 Derivative expansion of the pressure
${\cal L}=$ $P(\mu)-\frac{\partial P}{\partial\mu}v_{\mu}\partial^{\mu}\tilde{\theta}$
$+\frac{1}{2}\left[\eta^{\mu\nu}\frac{1}{\mu}\frac{\partial P}{\partial\mu}+v^{\mu}v^{\nu}\left(\frac{\partial^{2}P}{\partial\mu^{2}}-\frac{1}{\mu}\frac{\partial P}{\partial\mu}\right)\right]\partial_{\nu}\tilde{\theta}\partial_{\mu}\tilde{\theta}\,,$ (184)